OpenStudy (anonymous):

two circle A and B touch each other externally,PM=15 cm is a tangent to circle with centre Aand QN=13cm is a tangent to a circle with centre B from external points P and Q .if PA=17cm and BQ=12 cm . find the distance between the centres A and B of circles

OpenStudy (anonymous):



OpenStudy (anonymous):

ITS AN EASY ONE SOLVE IT COMMON GUYS

OpenStudy (anonymous):

@ikram

OpenStudy (anonymous):

@ganeshie8

OpenStudy (***[isuru]***):



OpenStudy (***[isuru]***):

the distance between centers of 2 circles = AC + CB but AC = AM (radius) CB = BN (radius) so the distance between the centers of 2 circles = AM + BN to find AM use Pythagoras theorem on triangle APM to find BN use Pythagoras theorem on triangles BNC can u do that ?

OpenStudy (***[isuru]***):

PM^2 + AM^2 = AP^2 AM^2 = AP^2 - PM^2 = 17^2 - 15^2

OpenStudy (anonymous):

or BC =BN AND AC=AM (radii of the circle)

OpenStudy (***[isuru]***):

there is something wrong with ur question BC should be larger than NC

OpenStudy (anonymous):

no in exm they had put it like this only

OpenStudy (***[isuru]***):

angle BNQ = 90 ( NQ is a tangent ) so triangle BNQ is a right angle triangle which means BQ is the hypotenuse... hypotenuse is the longest side in a triangle but ur data does not match it ???

OpenStudy (anonymous):

AC = 8 cm

OpenStudy (anonymous):

BC =5 cm

OpenStudy (***[isuru]***):

BC = 5cm if and only if BQ = 13 cm NQ = 12 cm not the other way around

OpenStudy (ikram002p):

u oly need to apply Pythagoras nothing more :o